Given:
A=
4 0 0
2 2 3
1 1 1
B=
4 0 0
2 2 3
1 1 1

Find: a) –A-1+ 3BT b) B-1+ ( AT+A-1)
1
Expert's answer
2020-04-30T15:18:14-0400

a) -A-1+3BT

"B^T=\begin{pmatrix}
 4 & 2 & 1 \\
 0 & 2 & 1 \\
 0 & 3 & 1
\end{pmatrix}"

"A^{-1}=\frac {1}{-4} \begin{pmatrix}
 -1 & 0 & 0 \\
 1 & 4 & -12 \\
 0 & -4 & 8
\end{pmatrix}=\begin{pmatrix}
 0.25 & 0 & 0 \\
 -0.25 & -1 & 3 \\
0 & 1 & -2
\end{pmatrix}"

"-A^{-1}+3B^T=\begin{pmatrix}
 -0.25 & 0 & 0 \\
 0.25 & 1 & -3 \\
0 & -1 & 2
\end{pmatrix}+\begin{pmatrix}
 12 & 6 & 3 \\
 0 & 6 & 3 \\
0 & 9 & 3
\end{pmatrix}=\begin{pmatrix}
 11.75 & 6 & 3 \\
 0.25 & 7 & 0 \\
0 & 8 & 5
\end{pmatrix}"

Answer:

"-A^{-1}+3B^T=\begin{pmatrix}
 11.75 & 6 & 3 \\
 0.25 & 7 & 0 \\
0 & 8 & 5
\end{pmatrix}"


b) B-1+(AT+A-1)

"B^{-1}+(A^T+A^{-1})=\begin{pmatrix}
 0.25 & 0 & 0 \\
 -0.25 & -1 & 3 \\
0 & 1 & -2
\end{pmatrix}+(
\begin{pmatrix}
 4 & 2 & 1 \\
 0 & 2 & 1 \\
 0 & 3 & 1
\end{pmatrix}+"

"\begin{pmatrix}
 0.25 & 0 & 0 \\
 -0.25 & -1 & 3 \\
0 & 1 & -2
\end{pmatrix})=\begin{pmatrix}
 4.5 & 2 & 1 \\
 -0.5 & 0 & 7 \\
0 & 5 & -3
\end{pmatrix}"

Answer:

"B^{-1}+(A^T+A^{-1})=\begin{pmatrix}
 4.5 & 2 & 1 \\
 -0.5 & 0 & 7 \\
0 & 5 & -3
\end{pmatrix}"